OpenStudy (anonymous):

What is a quartic function with only the two real zeros given? x = -4 and x = -1 A. y = x4 + 5x3 + 5x2 + 5x + 4 B. y = x4 - 5x3 - 5x2 - 5x - 4 C. y = -x4 + 5x3 + 5x2 + 5x + 4 D. y = x4 + 5x3 + 5x2 + 5x - 5

OpenStudy (anonymous):

@Forever_Me @rajee_sam

OpenStudy (anonymous):

For this, the factors must be (x+1)(x+4) to make it equal zero at both indicated values.

OpenStudy (anonymous):

I didn't get it

OpenStudy (anonymous):

@Numb3r1

OpenStudy (rajee_sam):

he is asking for Quartic function it has only 2 real zeros

OpenStudy (anonymous):

To make x=-4 and x=-1 give zero when fed through the function, (x+1)(x+4) must be the factorization of the equation. There must also be either a quadratic equation without any real zeros or more of those same factors.

OpenStudy (rajee_sam):

So other 2 are complex

OpenStudy (rajee_sam):

All I can think of is use synthetic division to divide all the 4 answer choices to see which one has both -1 and -4 as its zeros and then examine the quotient to see which one has complex roots

OpenStudy (anonymous):

he is asking which eqution has only the givin zeros -1 -4

OpenStudy (anonymous):

It isn't too hard to multiply those factors, then use synthetic division to test that. It'll save a little bit of work that way.

OpenStudy (rajee_sam):

It is very laborious but I cannot think of any other way now.

OpenStudy (anonymous):

I'd divide by the quadratic, is what I'm saying. It saves a step of division.

OpenStudy (anonymous):

guys I really don't know how to do this so can u help me figure out what is the answer

OpenStudy (rajee_sam):

I cannot do synthetic then, I would have to do long division once I multiply them.

OpenStudy (rajee_sam):

Ok do you know how to do synthetic division?

OpenStudy (anonymous):

Why not? It's one division instead of two.

OpenStudy (anonymous):

I'm really confused

OpenStudy (rajee_sam):

Once you multiply the two factors it will be a trinomial. Can I do synthetic with a trinomial divisor? We will help you out malka in a bit just hang on. Thinking of a less tedious way to do it.

OpenStudy (rajee_sam):

you would have to do the long division

OpenStudy (rajee_sam):

Ok here is what we will do. If x=-4 and x=-1 are zeroes. Then x+4 and x+1 are factors . Do you agree @malkassir

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So now multiply those.

OpenStudy (rajee_sam):

Now let us multiply the two together : (x+1)(x+4) = x^2 + 5x + 4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now we're going to do long division:

OpenStudy (rajee_sam):

Now take that and divide all the 4 answer choices. Once you divide it should divide completely without a remainder because they are a factor

OpenStudy (anonymous):

Here are your original options: A. y = x4 + 5x3 + 5x2 + 5x + 4 B. y = x4 - 5x3 - 5x2 - 5x - 4 C. y = -x4 + 5x3 + 5x2 + 5x + 4 D. y = x4 + 5x3 + 5x2 + 5x - 5

OpenStudy (rajee_sam):

@Numb3r1 , I will do A and show him how to do it.

OpenStudy (anonymous):

okay I will do B and @Numb3r1 you do c and we will see which one is right

OpenStudy (anonymous):

Here's a drawing to start all of them off:

OpenStudy (rajee_sam):